# The Impossibility of Inverse Permutation Learning in Transformer Models

### v1

@@ 2 Statement of Technical Results
We now turn to a more surprising result — simply padding the input is also sufficient to demonstrate a construction that solves the inverse permutation learning task. Let $s\in\mathbb{R}^{d}$ be the embedding associated with a special scratch token. Let $S=\mathbf{1}s^{\top}\in\mathbb{R}^{d\times d}$ be the matrix with $d$ rows where each row is $s$. The content of the representation does not matter for our purposes, so it may overlap with a regular token embedding or be an embedding not corresponding to any regular token. We state this theorem below.
@@
We provide a proof in Appendix˜C. The proof highlights that the specific content of the “padding” tokens (here, the $S$) is unimportant. Instead, as discussed in Section˜1, these padding tokens provide the model with additional “scratch space” to perform the necessary matrix operations. Unlike Theorem˜2, this result provides a recipe for performing inverse permutation learning without modifying the decoder-only architecture which underpins most modern LLMs.
@@ Appendix B Empirical validation
After training the model without the causal mask, the MSE of the model is about 0.00015. We reproduce the weights corresponding to the trained model in Figure˜2. Panel (a) shows the weights for the first attention layer $A^{(1)}$, panel (b) is for the second attention layer $A^{(2)}$ and (c) is for the output layer $W$. The construction recovered by the model training process is different than the construction provided in our proof of Theorem˜2; both are valid. We visualize weights with a heatmap: each small square in each figure represents a single weight parameter. Thus, $A^{(1)}$ is a matrix of size $3d\times 3d=30\times 30$. In (a), we label weights corresponding to token and position embeddings above and to the side of the matrix. Weights that are larger are closer to yellow and weights that are close to zero are dark blue. Thus, the weights matrix $A^{(1)}$ is approximately equal to
@@
Figure 2: Weights for decoder-only disentangled transformer trained to inverse permutations.
@@ C.2 Proof of Theorem 2
We proceed by construction. The first attention layer computes the row-reversed column embedding matrix $P$. For constant $\beta_{1}$, we will have attention weights represented by the following block matrix
${A}^{(1)}=\beta_{1}\left[{\begin{matrix}0&0&0\\ 0&J_{d}&0\\ 0&0&0\end{matrix}}\right]$
where each 0 is the $d\times d$ all-zeros matrix and $J_{d}\in\mathbb{R}^{d\times d}$ is the antidiagonal matrix with $(J_{d})_{i,j}=1$ if $i+j=d+1$ and 0 otherwise. This implies pre-activations
$X{A}^{(1)}X^{\top}$ $=X\left[{\begin{matrix}0&0\\ J_{d}&0\\ 0&0\end{matrix}}\right]$
$=\beta_{1}\left[{\begin{matrix}J_{d}&0\\ 0&0\end{matrix}}\right]\in\mathbb{R}^{T\times T}$
As $\beta_{1}\to\infty$, this implies activations
$\mathcal{S}(X{A}^{(1)}X^{\top})=\left[{\begin{matrix}J_{d}&0\\ 0&0\end{matrix}}\right]$
Finally, we have layer-1 outputs
$\mathcal{S}(X{A}^{(1)}X^{\top})X=\left[{\begin{matrix}P^{\prime}&J_{d}&0\\ 0&0&0\end{matrix}}\right]$
where $P^{\prime}$ is the matrix such that $P^{\prime}_{i,j}$ is $P_{(d+1-i),j}$ (i.e., the row-reversed matrix).
The second attention layer computes the column-reversed transformation of $P^{\prime}$, which is $P^{\top}$. Recall that the input to the second attention is
@@
$={\begin{bmatrix}P&I_{d}&0&P^{\prime}&J_{d}&0\\ Y_{P}&0&I_{d}&0&0&0\end{bmatrix}}$
@@
${A}^{(2)}=\beta_{2}\left[{\begin{array}[]{ccc|ccc}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \hline\cr 0&0&J_{d}&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \end{array}}\right]$
@@
$h^{(1)}{A}^{(2)}h^{(1)\top}$ $=\beta_{2}h^{(1)}\left[{\begin{matrix}0&0\\ 0&0\\ 0&0\\ 0&J_{d}\\ 0&0\\ 0&0\\ \end{matrix}}\right]$
$=\beta_{2}\left[{\begin{matrix}0&P^{\top}\\ 0&0\end{matrix}}\right]$
and, as $\beta_{2}\to\infty$, activations
$\mathcal{S}(h^{(1)}{A}^{(2)}h^{(1)\top})$ $=\left[{\begin{matrix}0&P^{\top}\\ 0&0\end{matrix}}\right]$
@@
$\mathcal{S}(h^{(1)}{A}^{(2)}h^{(1)\top})h^{(1)}={\begin{bmatrix}P^{\top}Y_{P}&0&P^{\top}&0&0&0\\ 0&0&0&0&0&0\end{bmatrix}}$
Finally, notice $P^{\top}Y_{P}=Y$ so the top-left block recovers the desired canonical order. ∎
@@ C.3 Proof of Theorem 3
Observe
@@
$=\left[{\begin{matrix}\begin{matrix}P\\ Y_{P}\\ S\end{matrix}&I_{T}\\ \end{matrix}}\right]$
$=\left[{\begin{matrix}P&I_{d}&0_{d}&0_{d}\\ Y_{P}&0_{d}&I_{d}&0_{d}\\ S&0_{d}&0_{d}&I_{d}\\ \end{matrix}}\right]\in\mathbb{R}^{T\times(d+T)}$
The first layer produces the row-reversed matrix $P^{\prime}$. For constant $\beta_{1}$, we will have attention weights represented by the following block matrix
${A}^{(1)}=\beta_{1}\left[{\begin{matrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&J_{d}&0&0\end{matrix}}\right]$
where each 0 is the $d\times d$ all-zeros matrix and $J_{d}\in\mathbb{R}^{d\times d}$ is the antidiagonal matrix with $(J_{d})_{i,j}=1$ if $i+j=d+1$ and 0 otherwise. This implies pre-activations
$X{A}^{(1)}X^{\top}$ $=\beta_{1}X\left[{\begin{matrix}0&0&0\\ 0&0&0\\ 0&0&0\\ J_{d}&0&0\end{matrix}}\right]$
$=\beta_{1}\left[{\begin{matrix}0&0&0\\ 0&0&0\\ J_{d}&0&0\end{matrix}}\right]\in\mathbb{R}^{T\times T}$
And as $\beta_{1}\to\infty$,
$\mathcal{S}(\mathrm{MASK}(X{A}^{(1)}X^{\top}))$ $=\begin{bmatrix}0&0&0\\ 0&0&0\\ J_{d}&0&0\end{bmatrix}$
This yields layer 1 outputs:
$\mathcal{S}(\mathrm{MASK}(X{A}^{(1)}X^{\top}))X$ $={\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ P^{\prime}&J_{d}&0&0\end{bmatrix}}\in\mathbb{R}^{T\times(d+T)}$
@@
$h^{(1)}$ $=\begin{bmatrix}P&I_{d}&0_{d}&0_{d}&0_{d}&0_{d}&0_{d}&0_{d}\\ Y_{P}&0_{d}&I_{d}&0_{d}&0_{d}&0_{d}&0_{d}&0_{d}\\ S&0_{d}&0_{d}&I_{d}&P^{\prime}&J_{d}&0_{d}&0_{d}\end{bmatrix}$
@@
${A}^{(2)}=\beta_{2}\left[{\begin{array}[]{cccc|cccc}0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ \hline\cr 0&0&J_{d}&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{array}}\right]$
@@
$h^{(1)}A^{(2)}h^{(1)\top}$ $=\beta_{2}h^{(1)}\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\\ 0&0&0\\ 0&J_{d}&0\\ 0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}$
$=\beta_{2}\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&P^{\top}&0\end{bmatrix}$
and activations
$\mathcal{S}(\mathrm{MASK}(h^{(1)}A^{(2)}h^{(1)\top}))$ $=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&P^{\top}&0\end{bmatrix}$.
@@
$\mathcal{S}(\mathrm{MASK}(h^{(1)}A^{(2)}h^{(1)\top}))h^{(1)}$ $=\begin{bmatrix}0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ Y&0&P^{\top}&0&0&0&0&0\\ \end{bmatrix}$

### v3

@@ 2 Statement of Technical Results
We now turn to a more surprising result — simply padding the input is also sufficient to demonstrate a construction that solves the inverse permutation learning task. Let $s\in\mathbb{R}^{d}$ be the embedding associated with a special scratch token. Let $S=\mathbf{1}s^{\top}\in\mathbb{R}^{d\times d}$ be the matrix with $d$ rows where each row is $s$. We will let $s=0_{1\times d}$ by convention. We state this theorem below.
@@
We provide a proof in Appendix˜C. The proof requires that the first token of the input be a special BOS token with both the token and positional embedding set to all zeros. The proof highlights that the specific content of the “padding” tokens (here, the $S$) is unimportant. Instead, as discussed in Section˜1, these padding tokens provide the model with additional “scratch space” to perform the necessary matrix operations. Unlike Theorem˜2, this result provides a recipe for inverse permutation learning without modifying the decoder-only architecture which underpins most modern LLMs.
@@ Appendix B Empirical validation
After training the model without the causal mask, the MSE of the model is about 0.00015. We reproduce the weights corresponding to the trained model in Figure˜2. Panel (a) shows the weights for the first attention layer $A^{(1)}$, panel (b) is for the second attention layer $A^{(2)}$ and (c) is for the output layer $W$. The construction recovered by the model training process is not the same as that of the construction provided in our proof of Theorem˜2; both are valid. We visualize weights with a heatmap: each small square in each figure represents a single weight parameter. Thus, $A^{(1)}$ is a matrix of size $3d\times 3d=30\times 30$. In (a), we label weights corresponding to token and position embeddings above and to the side of the matrix. Weights that are larger are closer to yellow and weights that are close to zero are dark blue. Thus, the weights matrix $A^{(1)}$ is approximately equal to
@@
Figure 2: Weights for decoder-only causal-mask-free disentangled transformer trained to invert permutations.
@@
Transformers with scratch space.
To facilitate ease of visual inspection of the trained transformers, we modify the training setup from above so that there is a $d$-dimensional all-zeroes matrix prepended to the input. After training the model, the test MSE is 0.00013. We reproduce the weights corresponding to the trained model in Figure˜3. The panels display the same quantities as in the case of the causal-mask-free transformer. These weights constitute an alternative constructive proof of Theorem 3.
(a) Weights for first attention layer $A^{(1)}$
(b) Weights for second attention layer $A^{(2)}$
(c) Weights for output layer $W$
Figure 3: Weights for decoder-only disentangled transformer with scratch space trained to invert permutations.
@@ C.2 Proof of Theorem 2
We proceed by construction.
The first layer copies $P$ into the bottom block-row of the residual stream. For constant $\beta_{1}$, we will have attention weights represented by the following block matrix
$A^{(1)}=\beta_{1}\left[{\begin{matrix}0_{d}&0_{d}&0_{d}\\ 0_{d}&0_{d}&0_{d}\\ 0_{d}&I_{d}&0_{d}\end{matrix}}\right]$
where each $0_{d}$ is the $d\times d$ all-zeros matrix and $I_{d}\in\mathbb{R}^{d\times d}$ is the $d\times d$ identity matrix. This implies pre-activations
$X{A}^{(1)}X^{\top}$ $=\beta_{1}\left[{\begin{matrix}0_{d}&0_{d}\\ I_{d}&0_{d}\end{matrix}}\right]\in\mathbb{R}^{T\times T}$
This implies activations
$\mathcal{S}(X{A}^{(1)}X^{\top})=\left[{\begin{matrix}1/(2d)\mathbf{1}\mathbf{1}^{\top}&1/(2d)\mathbf{1}_{d}\mathbf{1}_{d}^{\top}\\ I_{d}&0_{d}\\ \end{matrix}}\right]$
where $1/(2d)\mathbf{1}\mathbf{1}^{\top}$ is the $d\times d$ constant matrix with entries $1/(2d)$. Finally, we have layer-1 outputs
$\mathcal{S}(X{A}^{(1)}X^{\top})X=\left[{\begin{matrix}1/(2d)\mathbf{1}\mathbf{1}^{\top}(P+Y_{P})&\dots&\dots\\ P&\dots&\dots\end{matrix}}\right]$
where we write $\dots$ for block matrices that will not factor into later computations and therefore do not matter.
The second attention layer computes $Y$. Recall that the input to the second attention is
@@
$={\begin{bmatrix}P&I_{d}&0&1/(2d)\mathbf{1}\mathbf{1}^{\top}(P+Y_{P})&\dots&\dots\\ Y_{P}&0&I_{d}&P&\dots&\dots\end{bmatrix}}$
@@
${A}^{(2)}=\beta_{2}\left[{\begin{array}[]{ccc|ccc}0&0&0&0&0&0\\ 0&0&0&I&0&0\\ 0&0&0&0&0&0\\ \hline\cr 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \end{array}}\right]$
@@
$h^{(1)}{A}^{(2)}h^{(1)\top}$ $=\beta_{2}\left[{\begin{matrix}I_{d}\\ 0_{d}\end{matrix}}\right]\begin{bmatrix}1/(2d)\mathbf{1}\mathbf{1}^{\top}(P+Y_{P})^{\top}&P^{\top}\end{bmatrix}$
$=\beta_{2}\left[{\begin{matrix}1/(2d)\mathbf{1}\mathbf{1}^{\top}(P+Y_{P})^{\top}&P^{\top}\\ 0&0\end{matrix}}\right]$
Note that $1/(2d)\mathbf{1}\mathbf{1}^{\top}(P+Y_{P})^{\top}<1$ so, as $\beta_{2}\to\infty$, we have activations
$\mathcal{S}(h^{(1)}{A}^{(2)}h^{(1)\top})$ $=\left[{\begin{matrix}0&P^{\top}\\ 1/(2d)\mathbf{1}\mathbf{1}^{\top}&1/(2d)\mathbf{1}\mathbf{1}^{\top}\end{matrix}}\right]$
@@
$\mathcal{S}(h^{(1)}{A}^{(2)}h^{(1)\top})h^{(1)}={\begin{bmatrix}\dots&\dots&\dots&\dots&\dots&\dots\\ P^{\top}Y_{P}&\dots&\dots&\dots&\dots&\dots\end{bmatrix}}$
Finally, notice $P^{\top}Y_{P}=Y$ so that block recovers the desired canonical order. ∎
@@ C.3 Proof of Theorem 3
Observe, since we pad the beginning of the input with a BOS token to be all zeros, that
@@
$=\left[{\begin{matrix}0&0_{1\times T}\\ \begin{matrix}P\\ Y_{P}\\ S\end{matrix}&I_{T}\\ \end{matrix}}\right]$
$=\left[{\begin{matrix}0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}\\ P&I_{d}&0_{d}&0_{d}\\ Y_{P}&0_{d}&I_{d}&0_{d}\\ S&0_{d}&0_{d}&I_{d}\\ \end{matrix}}\right]\in\mathbb{R}^{(T+1)\times(d+T)}$
The first layer copies $P$ into the third block-row of the residual stream. For constant $\beta_{1}$, we will have attention weights represented by the following block matrix
${A}^{(1)}=\beta_{1}\left[{\begin{matrix}0&0&0&0\\ 0&0&0&0\\ 0&I_{d}&0&0\\ 0&0&0&0\end{matrix}}\right]$
where each 0 is the $d\times d$ all-zeros matrix. This implies pre-activations
$X{A}^{(1)}X^{\top}$ $=\beta_{1}\left[{\begin{matrix}0&0_{1\times d}&0_{1\times d}&0_{1\times d}\\ 0_{d\times 1}&0_{d}&0_{d}&0_{d}\\ 0_{d\times 1}&I_{d}&0_{d}&0_{d}\\ 0_{d\times 1}&0_{d}&0_{d}&0_{d}\end{matrix}}\right]\in\mathbb{R}^{(T+1)\times(T+1)}$
Let $\Delta^{(1)}$ be the lower triangular matrix with $\Delta_{ij}=1/i$ for $j\leq i$. Let $\Delta^{(1)}$ be the $d+1\times d+1$ matrix with $\Delta_{ij}=1/(2d+i+1)$. And as $\beta_{1}\to\infty$,
$\mathcal{S}(\mathrm{MASK}(X{A}^{(1)}X^{\top}))$ $=\begin{bmatrix}\Delta^{(1)}&&0&0\\ 0_{d\times 1}&I_{d}&0&0\\ 0_{d\times 1}&\Delta^{(2)}&\cdots&\cdots\end{bmatrix}$
We will write dots when the values in a block do not matter for subsequent computations. This yields layer 1 outputs:
$\mathcal{S}(\mathrm{MASK}(X{A}^{(1)}X^{\top}))X$ $={\begin{bmatrix}\Delta^{(1)}[0_{1\times d};P]&0_{1\times d}&0_{1\times d}&0_{1\times d}\\ \cdots&\cdots&\cdots&\cdots\\ P&0_{d\times d}&0_{d\times d}&0_{d\times d}\\ 0_{d}&\cdots&\cdots&\cdots\end{bmatrix}}\in\mathbb{R}^{T\times(d+T)}$
@@
$h^{(1)}$ $=\begin{bmatrix}0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}&\Delta^{(1)}[0_{1\times d};P]&0_{1\times d}&0_{1\times d}&0_{1\times d}\\ P&I_{d}&0_{d}&0_{d}&\cdots&\cdots&\cdots&\cdots\\ Y_{P}&0_{d}&I_{d}&0_{d}&P&0_{d\times d}&0_{d\times d}&0_{d\times d}\\ 0_{d}&0_{d}&0_{d}&I_{d}&0_{d}&\cdots&\cdots&\cdots\end{bmatrix}$
@@
${A}^{(2)}=\beta_{2}\left[{\begin{array}[]{cccc|cccc}0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&I_{d}&0&0&0\\ \hline\cr 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{array}}\right]$
@@
$h^{(1)}A^{(2)}h^{(1)\top}$ $=\beta_{2}\begin{bmatrix}0_{1\times d}\\ 0_{d}\\ 0_{d}\\ I_{d}\end{bmatrix}\begin{bmatrix}0_{d\times 1}&\Delta^{(1)}[0_{1\times d};P]&P&0_{d}\end{bmatrix}$
$=\beta_{2}\begin{bmatrix}0&0_{1\times d}&0_{1\times d}&0_{1\times d}\\ 0_{d\times 1}&0_{d}&0_{d}&0_{d}\\ 0_{d\times 1}&0_{d}&0_{d}&0_{d}\\ 0_{d\times 1}&\Delta^{(1)}[0_{1\times d};P]&P^{\top}&0_{d}\end{bmatrix}$.
Now, notice that all entries of $\Delta^{(1)}[0_{1\times d};P]$ are less than 1, since all non-zero entries of $P$ are 1 and all entries below the first row of $\Delta^{(1)}$ are less than 1. Thus, as $\beta_{2}\to\infty$, we have activations
$\mathcal{S}(\mathrm{MASK}(h^{(1)}A^{(2)}h^{(1)\top}))$ $=\begin{bmatrix}\dots&0&0\\ \dots&\dots&0\\ 0&P^{\top}&0\end{bmatrix}$.
@@
$\mathcal{S}(\mathrm{MASK}(h^{(1)}A^{(2)}h^{(1)\top}))h^{(1)}$ $=\begin{bmatrix}0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}&0_{1\times d}\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ Y&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ \end{bmatrix}$